3 white chips, 7 blue chips, 16 green chips, 2 chips drawn from the box in succession what is the probability that one is blue and other is white?

Answer by stanbon(75887)

3 white chips, 7 blue chips, 16 green chips, 2 chips drawn from the box in succession what is the probability that one is blue and other is white?
----------------------
# of ways to succeed: 7*3 = 21
----
# of possible outcomes:26C2 = (26*25)/(1*2) = 325
-----------
Ans: 21/325 = 0.0646
========================
Cheers,
Stan H.
========================